# Creating Multi-Level Skill Hierarchies in Reinforcement Learning

### v1

@@ 5 Empirical Analysis
We analyse the Louvain skill hierarchy in six environments: Rooms, Grid, Maze (Ramesh et al., 2019), Office, Taxi Dietterich (2000), and Towers of Hanoi. The environments are depicted in Figure 1 and fully described in Appendix D. In all environments, the agent receives a reward of $-0.001$ for each action and an additional $+1.0$ for reaching a goal state. In Rooms, Grid, Maze, and Office, there are 4 primitive actions: north, south, east, and west. In Taxi, there are two additional primitive actions: pick-up-passenger and put-down-passenger. Taxi also features irreversible actions. For instance, after picking up the passenger, the agent cannot return to a state where the passenger has not yet been picked up. All experimental details are in Appendix F.
Our analysis is directed by the following questions: What is the Louvain skill hierarchy produced in each environment? How does this skill hierarchy impact the learning performance of the agent? How are the results impacted as the environment gets larger? We report results using resolution parameter $\rho=0.05$ (unless stated otherwise) but also examine the impact of $\rho$ on the results.
@@
Figure 2: The cluster hierarchies produced by the Louvain algorithm when applied to the state-transition graphs representing Rooms, Office, Taxi, and Towers of Hanoi. For Taxi and Towers of Hanoi, the graph layout was determined by using a force-directed algorithm that models nodes as charged particles that repel each other and edges as springs that attract connected nodes.
Louvain Skill Hierarchy. Figure 2 shows the Louvain cluster hierarchy in Rooms, Office, Taxi, and Towers of Hanoi obtained by applying the Louvain algorithm to the state-transition graph of each environment. Appendix E shows the Louvain cluster hierarchies in Grid and Maze.
In Rooms, the hierarchy has four levels. At the third level, each room is placed in its own cluster. Moving up the hierarchy, at the fourth level, two of these rooms are joined together into a single cluster. Moving down the hierarchy, each room is divided further into smaller clusters at level 2, and then into even smaller clusters at level 1. It is easy to see how the corresponding skill hierarchy would enable efficient navigation between and within rooms.
In Office, at the top level, we see six large clusters connected to each other by corridors. As we move lower down the hierarchy, these clusters are divided into increasingly smaller regions. At level 3, we see many rooms that form their own cluster. At level 2, most rooms are divided into multiple clusters. Once again, it is relatively easy to see how the corresponding skill hierarchy would enable efficient navigation of the space at multiple levels of granularity.
In Taxi, the state-transition graph has four disconnected components, each corresponding to one particular passenger destination: R, G, B, or Y. In Figure 2, we show only one of these components, the one where the passenger destination is B. The Louvain hierarchy has four levels. At the top level, we see three clusters where the passenger is waiting at R, G, or Y, and a fourth cluster where the passenger is either in-taxi or delivered to the destination. Navigation between these clusters is unidirectional, with only three possibilities, and the three corresponding skills navigate the taxi to the passenger location and pick up the passenger. Moving one level down the hierarchy, the clusters correspond to skills that move the taxi between the left and the right side of the grid, which are connected by a bottleneck state in the middle of the grid.
In Towers of Hanoi, the hierarchy has three levels. Moving between the top-level, middle-level, and bottom-level clusters corresponds to moving, respectively, the largest, the second-largest, and the third-largest disc between the different poles.
In each domain, (1) the Louvain skill hierarchy closely matches human intuition, and (2) it is clear how skills at one level of the hierarchy can be composed to produce the skills at the next level.
@@
Figure 4: Learning curves comparing various different Louvain agents. An epoch corresponds to 100 decision stages in Rooms, 300 in Taxi, and 750 in Maze.
@@
Learning Performance. We compare learning performance with the Louvain skill hierarchy to the Edge Betweenness (Davoodabadi and Beigy, 2011) and Label Propagation (Farahani and Mozayani, 2019) methods, and to the method proposed by Xu et al. (Xu et al., 2018). These methods are the most directly related to the proposed approach because they make use of modularity maximisation to define their two-level skill hierarchies. In addition, we compare to options that navigate to local maxima of Node Betweenness Şimşek and Barto (2009), a state-of-the-art subgoal-based approach that captures the bottleneck concept that characterises many existing subgoal-based methods. We also compare to Eigenoptions Machado et al. (2017) derived from the graph Laplacian, a state-of-the-art Laplacian-based approach that encourages efficient exploration of the state space. Finally, we also include a Primitive agent that uses only primitive actions. Primitive actions are available to all agents.
For all methods, we generated options using the complete state-transition graph and learned their policies offline using macro-Q learning McGovern et al. (1997). We trained all hierarchical agents using macro-Q learning and intra-option learning Sutton and Precup (1998). Although these algorithms have not previously been applied to multi-level hierarchies, they both extend naturally to this case. The primitive agent was trained using Q-Learning Watkins (1989). The shaded regions on the learning curves represent the standard error over $40$ random seeds. When creating the Louvain skill hierarchy, we discarded any lower level of the cluster hierarchy that had a mean number of nodes per cluster of less than a threshold value, $c$. We used $c=4$ in all experiments; our reasoning is that skills that navigate between such small clusters execute for only a very small number of decision stages (often only one or two) and are not meaningfully more abstract than primitive actions.
@@
We show learning curves in Figure 3. The Louvain agent has a clear and substantial advantage over other approaches in all domains except for Towers of Hanoi, where its performance was much closer to that of the other hierarchical agents, with none of them performing much better than the primitive agent. This is consistent with existing results reported in this domain (e.g., by Jinnai et al. (2019)).
Hierarchical versus flat arrangement of skills. An alternative to the Louvain skill hierarchy is a flat arrangement of the same skills, where the skills have the same behaviour but they call primitive actions directly rather than indirectly through other (lower-level) skills. We expect the multi-level hierarchy to lead to faster learning than the flat hierarchy due to the additional macro-Q and intra-option learning updates enabled by the hierarchical relationship between the skills. Figure 4 shows that this is indeed the case. In the figure, Louvain shows an agent that uses the Louvain skill hierarchy while Louvain flat shows an agent that uses the Louvain skills but where the skill policies call primitive actions directly rather than through other skills. In addition, the figure shows a number of agents that use only a single level of the Louvain hierarchy, with option policies defined over primitive actions; these are depicted in the figure by the label Level 1, 2, 3, or 4. Primitive actions were available to all agents. The figure shows that the hierarchical agent learns more quickly than the flat agent. Furthermore, the agents using individual levels of the Louvain hierarchy learn more quickly than the primitive agent but not as quickly as the agent using the full Louvain hierarchy.
@@
Figure 5: (a) How the Louvain algorithm’s output when applied to Towers of Hanoi changes with the resolution parameter. (b) How the Louvain skill hierarchy’s depth scales with the size of the state space. (c) Learning performance in Office with two floors containing $2537$ states.
Impact of the resolution parameter $\boldsymbol{\rho}$. Figure 4(a) shows how changing $\rho$ impacts the cluster hierarchy produced by the Louvain algorithm in Towers of Hanoi. At $\rho=10$, the output is a single level containing many small clusters comprised of three nodes. At $\rho=3.3$, a two-level cluster hierarchy is produced: level 1 is identical to the partition produced at $\rho=10$, but level 2 contains larger clusters, each formed by merging three of the clusters from level 1. Further decreasing $\rho$ produces additional levels, each containing progressively fewer, larger clusters. As $\rho$ is reduced, the clusters identified at a given level of the hierarchy generally remain stable: the lowest-level partition found using a higher value of $\rho$ will typically be the same as the lowest-level partition found using a lower value. In other words, decreasing $\rho$ may add levels to the hierarchy, but it generally does not impact the existing levels. A sensitivity analysis on the value of $\rho$ showed that a wide range of $\rho$ values led to useful skills being produced, and that performance gradually decreased to no worse than that of a primitive agent at higher values of $\rho$. Please refer to Appendix A for this sensitivity analysis and a more detailed discussion on the choice of $\rho$.
How do Louvain skills scale to larger domains? The Louvain algorithm has been successfully applied to graphs with millions of nodes and billions of edges in minutes Blondel et al. (2008), and has been observed empirically to have a time complexity that is linear in the number of graph edges Blondel et al. (2008); Lancichinetti and Fortunato (2009).
Also important is how the Louvain skill hierarchy changes with the size of the state space. We experimented with a multi-floor version of the Office environment, with floors connected by a central elevator, where two primitive actions move the agent up and down between adjacent floors. The size of the state space can be varied by adjusting parameters such as the number of office floors. We generated a series of fifteen offices of increasing size, with the smallest office having a single floor (${\sim}10^{3}$ states), and the largest office having one thousand floors (${\sim}10^{6}$ states). Figure 4(b) shows that hierarchy depth increased very gradually with the size of the state space. At $1000$ states, the hierarchy contained $5$ levels, with the top-level skills allowing high-level navigation of each office floor, as well as from regions near an elevator to adjacent floors. At ${\sim}4000$ states, a sixth level was added, containing skills for efficiently moving from anywhere on one floor to an adjacent floor. Subsequently, as the state space size grew to ${\sim}25000$ and ${\sim}300000$ states, two levels were added that allowed the agent to move, respectively, two and four office floors at a time. Figure 4(c) shows that the Louvain agent learns much more quickly than other approaches in a two-floor Office, even while some alternatives, including the Eigenoptions agent, fail to achieve any learning.
@@ 6 Discussion and Future Work
Our results show that the Louvain skill hierarchy is a useful answer to the conceptual question of what constitutes a useful skill hierarchy. This hierarchy is intuitively appealing, improves agent performance (more so than alternatives in the literature), and shows desirable scaling properties.
@@
Figure 6: (a) Performance of Incremental Louvain Options in Rooms. An epoch corresponds to $100$ decision stages. (b-d) How the state-transition graph and top-level partitions produced by the Incremental 2 method evolved as an agent explored Rooms. The hierarchy contained 2 levels after visiting 30 states, 3 levels after visiting 60 states, and 5 levels after observing all possible transitions.
An important research direction for future work is incremental learning of Louvain skill hierachies as the agent is interacting with its environment—the state-transition graph will not always be available in advance. We explored the feasibility of incremental learning in the Rooms environment and present the results in Figure 6. The agent started with an empty state-transition graph and no skills. Every $m$ decision stages, it updated its state-transition graph with new nodes and edges, and it revised its skill hierarchy in one of two ways. In the first approach, the agent applied the Louvain algorithm to create a new skill hierarchy from scratch. In the second approach, the agent incorporated the new information into its existing skill hierarchy, using an algorithm similar to the Louvain algorithm. This algorithms starts by assigning each new node to its own cluster; it is then iteratively moved locally, between neighbouring clusters (both new and existing), until no modularity gain is possible. This revised partition is used to define an aggregate graph and the entire process is then repeated on the aggregate graph, and the next aggregate graph, and so on, until an iteration is reached with no modularity gain. Aside from existing clusters being merged into new higher-level clusters, the cluster membership of existing nodes stays fixed; only new nodes have their cluster membership updated. The result is a revised set of partitions from which a revised hierarchy of Louvain skills are derived. Pseudocode for these incremental approaches is in Appendix C.
Figures 5(b)5(d) show the evolution of the partitions using the second approach as the agent performed a random walk in Rooms. The partitions were updated after observing $30$ states, $60$ states, and all possible transitions. The figure shows that, as more nodes were added, increasingly higher-level skills were produced. After $30$ states, the top-level skills allow low-level movement within and between two of the rooms. After $60$ states, another level was added, allowing slightly higher-level navigation of the state-space. After observing all possible transitions, the top level contained skills enabling efficient movement between the four rooms.
@@
Figure 5(a) shows the performance of the incremental agents. The agents started with only primitive actions; after decision stages $100$, $500$, $1000$, $3000$, and $5000$, the state-transition graph was updated and the skill hierarchy was revised following the two approaches discussed above. The figure compares performance to a Primitive agent and an agent using the full Louvain skill hierarchy, whose performance acts as a Ceiling for the incremental agents. Each incremental agent learned much faster than the primitive agent and only marginally slower than the fully-informed Louvain agent. The two incremental agents had similar performance throughout training but the first approach reached a higher level of asymptotic performance than the second approach. The reason is that partitions produced early in training are based on incomplete information; the first approach discards the early imperfections while the second approach carries them forward. But there is a trade-off: the first approach has a higher computational cost than the second one. These results demonstrate the feasibility of learning Louvain skills incrementally. A full incremental method for learning Louvain skills may take many forms, and different approaches may be useful under different circumstances, with each having its own advantages and disadvantages. We leave the full development of such algorithms to future work.
@@
Another direction for future work is extending Louvain skills to environments with continuous state spaces such as robotic control tasks. Such domains present a difficulty to all graph-based skill discovery methods due to the inherently discrete nature of the state-transition graph. If the critical step of constructing an appropriate graphical representation of a continuous state space can be achieved, all graph-based methods would benefit. Some approaches have already been proposed in the literature; we use one such approach Mahadevan and Maggioni (2007); Bacon and Precup (2013); Shoeleh and Asadpour (2017) to examine the Louvain hierarchy in a variant of the Pinball domain Konidaris and Barto (2009), which involves manoeuvring a ball around a series of obstacles to reach a goal, shown in Figure 6(a). The state is represented by two continuous values: the ball’s position in the $x$ and $y$ directions. At each decision stage, the agent can choose to apply a small force to the ball in each direction. The amount of force applied is stochastic and causes the ball to roll until friction causes it to come to a rest. Collisions between the ball and the obstacles are elastic. We sampled $4000$ states, added them to the state-transition graph, then added an edge between each node and its $k$-nearest neighbours, according to euclidean distance, assigning each edge $(u,v)$ a weight of $e^{-\nicefrac{{\lVert u-v\rVert^{2}}}{{\sigma}}}$. We then applied the Louvain algorithm to the resulting graph and derived the Louvain skill hierarchy from the partitions produced. The result was the three-level cluster hierarchy shown in Figure 7. The highest-level clusters yield skills for high-level navigation of the state space and take into account features such as the natural bottlenecks caused by the obstacles, allowing the agent to efficiently change its position. Skills derived from the lower-level partitions enable more local navigation. Once again, we see a skill hierarchy that is intuitive, and it can easily be seen how the lower-level skills can be composed to produce the higher-level skills in the hierarchy.
(a) Pinball Layout
(b) Level 3
(c) Level 2
(d) Level 1
@@
Figure 7: (a) Pinball domain, with the green ball in its initial position, the red goal, and several obstacles. (b–d) The cluster hierarchy produced by the Louvain algorithm.
@@
Currently, Louvain skills at one level of the hierarchy are composed of skills from only the immediately previous level. While such skills may be optimal with respect to the skills from the previous level, they may not be optimal with respect to primitive actions. Future work could consider higher-level skills composed of skills from all lower levels, including primitive actions.
@@
Because we derive Louvain skills solely from the connectivity of the state-transition graph, not considering the reward function, we expect them to be suitable for solving a range of tasks in a given domain. Examining their use for transfer learning is a useful direction for future work. Additionally, future work can examine how to use the reward function when defining Louvain skills to tailor the resulting skills for solving specific tasks.
@@
A possible difficulty with building multi-level skill hierarchies is that having a large number of skills can end up hurting performance by increasing the branching factor of the problem. Future work should consider how best to manage large skill hierarchies. One solution that has been explored in the context of two-level hierarchies is “pruning” less useful skills from the hierarchy Taghizadeh and Beigy (2013); Farahani and Mozayani (2020).
@@
Lastly, we point out that the various characterisations of useful skills proposed in the literature, including the one proposed here, are not necessarily competitors. To solve complex tasks, it is likely that an agent will need to use many different types of skills. An important avenue for future work is studying how different ideas on skills discovery can be brought together to enable agents that can autonomously develop complex and varied skill hierarchies.
@@
This research was supported by the Engineering and Physical Sciences Research Council [EP/R513155/1] and the University of Bath. This research made use of Hex, the GPU Cloud in the Department of Computer Science at the University of Bath. We would like to thank the members of the Bath Reinforcement Learning Laboratory for their constructive feedback.

### v2

@@ 5 Empirical Analysis
We analyse the Louvain skill hierarchy in the six environments depicted in Figure 1: Rooms, Grid, Maze (Ramesh et al., 2019), Office, Taxi Dietterich (2000), and Towers of Hanoi. In all environments, the reward is $-0.001$ for each action and an additional $+1.0$ for reaching a goal state. In Rooms, Grid, Maze, and Office, there are four primitive actions: north, south, east, and west. In Taxi, there are two additional primitive actions: pick-up-passenger and put-down-passenger. Some decisions in Taxi are irreversible. For instance, after picking up the passenger, the agent cannot return to a state where the passenger is still waiting to be picked up. We describe each of these environments fully in Section A of the supplementary material.
@@
When generating partitions of the state transition graph using the Louvain algorithm, we used a resolution parameter of $\rho=0.05$, unless stated otherwise. When converting the output of the Louvain algorithm into a concrete skill hierarchy, we discarded all levels of the cluster hierarchy where the mean number of nodes per cluster was less than $4$. Our reasoning is that skills that navigate between such small clusters execute for only a very small number of decision stages (often only one or two) and are not meaningfully more abstract than primitive actions. For all methods used in our comparisons, we generated options using the complete state transition graph and learned their policies offline using macro-Q learning McGovern et al. (1997). We trained all hierarchical agents using macro-Q learning and intra-option learning Sutton and Precup (1998). Although these algorithms have not previously been applied to multi-level hierarchies, they both extend naturally to this case. The primitive agent was trained using Q-Learning Watkins (1989). The shaded regions on the learning curves represent the standard error in $40$ independent runs. All experiments are fully described in Section B of the supplementary material.
Our analysis is directed by the following questions: What is the Louvain skill hierarchy generated in each environment? How does this skill hierarchy impact the learning performance of the agent? How do the results change as the number of states increases? Does arranging skills into a multi-level hierarchy provide benefits over a flat arrangement of the same skills? What is the impact of varying the value of the resolution parameter $\rho$?
Louvain Skill Hierarchy. We first examine the cluster hierarchies generated by applying the Louvain algorithm to the state transition graphs of various environments. Figure 2 shows the results in Rooms, Office, Taxi, and Towers of Hanoi. Section D of the supplementary material shows additional results in Grid and Maze.
@@
Figure 2: The cluster hierarchies produced by the Louvain algorithm when applied to the state transition graphs representing Rooms, Office, Taxi, and Towers of Hanoi. For Taxi and Towers of Hanoi, the graph layout was determined by using a force-directed algorithm that models nodes as charged particles that repel each other and edges as springs that attract connected nodes.
In Rooms, the hierarchy has four levels. At level three, we see that each room has been placed in its own cluster. Moving up the hierarchy, at level four, two of these rooms have been joined together into a single cluster. Moving down the hierarchy, each room has been divided into smaller clusters at level two, and then into even smaller clusters at level one. The figure illustrates how the corresponding skill hierarchy would enable efficient navigation between and within rooms.
In Office, at the top level, we see six large clusters connected by corridors. As we move lower down the hierarchy, we see that these large clusters have been divided into increasingly smaller regions. At level three, there are many rooms that form their own cluster. At level two, most rooms have been divided into multiple clusters. The figure reveals how the corresponding skill hierarchy would enable efficient navigation of the environment at multiple time scales.
In Taxi, the state transition graph has four disconnected components, each corresponding to one particular passenger destination: R, G, B, or Y. In Figure 2, we show only one of these components, the one where the passenger destination is B. The results for the other components are similar. The Louvain cluster hierarchy has four levels. At the top level, we see four clusters. In three of these clusters, the passenger is waiting at their starting location (R, G, or Y). In the fourth cluster, the passenger is either in-taxi or has been delivered to their destination (B). Navigation between these clusters is unidirectional, with only three possibilities. The three corresponding skills navigate the taxi to the passenger location and pick up the passenger. Moving one level down the hierarchy, the clusters produce skills that move the taxi between the left and the right-hand side of the grid, which are connected by the bottleneck state at the centre of the taxi navigation grid.
In Towers of Hanoi, the hierarchy has three levels. At level three, moving between the three clusters corresponds to moving the largest disk to a different pole. Each of these clusters has been divided into three smaller clusters at level two and then into three even smaller clusters at level one. A similar structure exists at levels two and one. The smaller clusters at level two correspond to moving the second-largest disc between different poles; the smallest clusters at level one correspond to moving the third-largest disc between different poles.
In all of these domains, the Louvain skill hierarchy closely matches human intuition. In addition, it is clear how skills at one level of the hierarchy can be composed to produce the skills at the next level.
@@
Learning Performance. We compare learning performance with the Louvain skill hierarchy to approaches based on Edge Betweenness (Davoodabadi and Beigy, 2011) and Label Propagation (Farahani and Mozayani, 2019), as well as to the method proposed by Xu et al. (Xu et al., 2018). These methods are the most directly related to the proposed approach, as discussed in Section 4. In addition, we compare to options that navigate to local maxima of Node Betweenness Şimşek and Barto (2009), a subgoal-based approach that captures the many different notions of the bottleneck concept in the literature. Similarly to the proposed approach, it is also one of the few approaches in the literature to explicitly characterise a target skill hierarchy for an agent to learn. We also compare to Eigenoptions Machado et al. (2017), which are derived from the Laplacian of the state transition graph and encourage efficient exploration. Finally, we include a Primitive agent that uses only primitive actions. Primitive actions are available to all agents.
@@
We present learning curves in Figure 3. The Louvain agent has a clear and substantial advantage over other approaches in all domains except for Towers of Hanoi, where its performance is much closer to that of the other hierarchical agents, with none of them performing much better than the primitive agent. This is consistent with existing results reported in this domain (e.g., by Jinnai et al. (2019)).
Section C of the supplementary material contains further analysis comparing Louvain skills and skills based on node betweenness.
Scaling to larger domains. We experimented with a multi-floor version of the Office environment, with floors connected by a central elevator, where two primitive actions move the agent up and down between two adjacent floors. The number of states in this domain can be varied by adjusting parameters such as the number of office floors. We use this environment to explore how the performance of the agent and the Louvain skill hierarchy change as the number of states in the environment increases.
@@
Figure 4: (a) Learning performance in a two-floor Office containing $2537$ states. (b) How the depth of the skill hierarchy scales with the size of the state space in multi-floor Office.
Figure 3(a) presents results in an Office environment with $2537$ states. It shows that the Louvain agent learns much more quickly than all other agents. Some alternatives, including the Eigenoptions agent, fail to achieve any learning.
Figure 3(b) shows how the Louvain skill hierarchy changes with the size of the state space in the Office environment. The figure shows the depth of the skill hierarchy in a series of fifteen Office environments of increasing size, ranging from a single floor (${\sim}10^{3}$ states) to one thousand floors (${\sim}10^{6}$ states). The depth of the hierarchy increased very gradually with the number of states in the environment. The maximum hierarchy depth reached was eight in the largest environment tested.
Figure 5: Learning curves comparing various different Louvain agents. An epoch corresponds to 100 decision stages in Rooms, 300 in Taxi, and 750 in Maze. The skill hierarchy contained three levels in Rooms and Taxi, and four levels in Maze.
Hierarchical versus flat arrangement of skills. An alternative to the Louvain skill hierarchy is a flat arrangement of the skills, where each skill calls primitive actions directly rather than indirectly through other (lower-level) skills. We expected the hierarchical arrangement to lead to faster learning than the flat arrangement due to the additional learning updates enabled by the hierarchical relationship between the skills. Figure 5 shows that this is indeed the case. In the figure, Louvain shows an agent that uses the Louvain skill hierarchy while Louvain flat shows an agent that uses the Louvain skills but where the skill policies call primitive actions directly rather than through other skills. In addition, the figure shows a number of agents that use only a single level of the Louvain hierarchy (Level 1, Level 2, Level 3, or Level 4), with option policies that call primitive actions directly. Primitive actions were available to all agents. The figure shows that the hierarchical agent learns more quickly than the flat agent. Furthermore, the agents using individual levels of the Louvain hierarchy learn more quickly than the primitive agent but not as quickly as the agent using the full Louvain hierarchy.
Figure 6: How the Louvain algorithm’s output when applied to Towers of Hanoi and Rooms changes with the resolution parameter. The cluster hierarchy had a maximum of four levels in Towers of Hanoi and five levels in Rooms.
Impact of the resolution parameter. When using very high values of the resolution parameter $\rho$, the Louvain algorithm terminates without producing any partitions. On the other hand, when using very low values of $\rho$, it runs until a partition is produced where all nodes are merged into a single cluster.
Figure 6 shows how changing the resolution parameter $\rho$ impacts the Louvain cluster hierarchy in Towers of Hanoi and Rooms. In Towers of Hanoi, at $\rho=10$, the output is a single level containing many small clusters comprised of three nodes. As $\rho$ gets smaller, the cluster hierarchy remains the same until $\rho=3.3$, at which point a second level is added. The first level remains identical to the single level produced at $\rho=10$. The second level contains larger clusters, each formed by merging three of the clusters from the first level. As $\rho$ is reduced further, additional levels are added to the hierarchy. When a new level is added to the hierarchy, the existing levels generally remain the same.
A sensitivity analysis on the value of $\rho$ showed that a wide range of $\rho$ values led to useful skills, and that performance gradually decreased to no worse than that of a primitive agent at higher values of $\rho$. Section E of the supplementary material contains full details of this sensitivity analysis and a more detailed discussion on the choice of $\rho$.
@@ 6 Discussion and Future Work
An important research direction for future work is incremental learning of Louvain skill hierachies as the agent interacts with its environment—because the state transition graph will not always be available in advance. We explored the feasibility of incremental learning in the Rooms environment. The results are shown in Figure 7. The agent started with an empty state transition graph and no skills. Every $m$ decision stages, it updated its state transition graph with new nodes and edges, and it revised its skill hierarchy using one of two approaches: Replace or Update.
In the first approach (Replace), the agent applied the Louvain algorithm to create a new skill hierarchy from scratch. In the second approach (Update), the agent incorporated the new information into its existing skill hierarchy, using an approach similar to the Louvain algorithm. This second approach starts by assigning each new node to its own cluster; the new nodes are then iteratively moved locally, between neighbouring clusters (both new and existing) until no modularity gain is possible. This revised partition is used to define an aggregate graph and the entire process is repeated on the aggregate graph, and the next aggregate graph, and so on, until an iteration is reached with no modularity gain. The cluster membership of existing nodes stays fixed; only new nodes have their cluster membership updated. The result of both approaches is a revised set of partitions, from which a revised hierarchy of Louvain skills is derived. Section I of the supplementary material presents pseudocode for the two incremental approaches.
@@
Figure 7: (a) Performance when incrementally discovering Louvain options in Rooms. An epoch corresponds to $100$ decision stages. (b-d) How the state transition graph and top-level partitions produced by the Update approach evolved as an agent explored Rooms. The hierarchy contained 2 levels after visiting 30 states, 3 levels after visiting 60 states, and 5 levels after observing all possible transitions.
@@
Figure 6(a) shows the performance of the incremental agents. The agents started with only primitive actions; after decision stages $100$, $500$, $1000$, $3000$, and $5000$, the state transition graph was updated and the skill hierarchy was revised. The figure compares performance to a Primitive agent and an agent using the full Louvain skill hierarchy, whose performance acts as a Ceiling for the incremental agents. Both incremental agents learned much faster than the primitive agent and only marginally slower than the fully-informed Louvain agent. The two incremental agents had similar performance throughout training but Replace reached a higher level of asymptotic performance than Update, as expected. The reason is that partitions produced early in training are based on incomplete information; Replace discards these early imperfections while Update carries them forward. But there is a trade-off: building a new skill hierarchy from scratch has a higher computational cost than updating an existing one.
Figures 6(b)6(d) show the evolution of the partitions as the Update agent performed a random walk in Rooms. The partitions were updated after observing $30$ states, $60$ states, and all possible transitions. The figure shows that, as more nodes were added, increasingly higher-level skills were produced.
These results demonstrate the feasibility of learning Louvain skills incrementally. A full incremental method for learning Louvain skills may take many forms, and different approaches may be useful under different circumstances, with each having its own strengths and weaknesses. We leave the full development of such algorithms to future work.
@@
Another direction for future work is extending Louvain skills to environments with continuous state spaces such as robotic control tasks. Such domains present a difficulty to all skill discovery methods that use the state transition graph due to the inherently discrete nature of the graph. If the critical step of constructing an appropriate graphical representation of a continuous state space can be achieved, all graph-based methods would benefit. Some approaches have already been proposed in the literature. We use one such approach Mahadevan and Maggioni (2007); Bacon and Precup (2013); Shoeleh and Asadpour (2017) to examine the Louvain hierarchy in a variant of the Pinball domain Konidaris and Barto (2009), which involves manoeuvring a ball around a series of obstacles to reach a goal, as shown in the leftmost panel of Figure 8. The state is represented by two continuous values: the ball’s position in the horizontal and vertical directions. At each decision stage, the agent can choose to apply a small force to the ball in each direction. The amount of force applied is stochastic and causes the ball to roll until friction causes it to come to a rest. Collisions between the ball and the obstacles are elastic.
Pinball Layout
Level 3
Level 2
Level 1
@@
Figure 8: The layout of the Pinball environment—with the green ball in its initial position, the red goal, and several obstacles—and the Louvain cluster hierarchy produced by the Louvain algorithm.
We sampled $4000$ states, added them to the state transition graph, then added an edge between each node and its $k$-nearest neighbours according to Euclidean distance, assigning an edge weight of $e^{-4d^{2}}$ to an edge that connects two locations $u$ and $v$ with Euclidean distance $d$ between them. We then applied the Louvain algorithm to the resulting graph. The result was the cluster hierarchy shown in Figure 8. It had three levels. Moving between clusters in the top level corresponds to skills that enable high-level navigation of the state space, and take into account features such as the natural bottlenecks caused by the obstacles, allowing the agent to efficiently change its position. Moving between clusters in the lower-level partitions enable more local navigation.
@@
Currently, Louvain skills at one level of the hierarchy are composed of skills from only the previous level. Such skills may not provide the most efficient navigation between clusters. Future work could consider composing skills from all lower levels, including primitive actions.
@@
Because Louvain skills are derived from the connectivity of the state transition graph, we expect them to be suitable for solving a range of problems in a given environment. Examining their use for transfer learning is a useful direction for future work.
@@
A possible difficulty with building multi-level skill hierarchies is that having a large number of skills can end up hurting performance by increasing the branching factor. One solution that has been explored in the context of two-level hierarchies is pruning less useful skills from the hierarchy Taghizadeh and Beigy (2013); Farahani and Mozayani (2020). Further research is needed to address this potential difficulty.
A key open problem for graph-based skill discovery methods generally is how to best make use of state abstraction. We suggest that the most natural place to introduce state abstraction is when constructing the state transition graph itself. Instead of a concrete state, each node could represent an abstract state based on some learned representation of the environment. The proposed method—and any other existing graph-based method—could then directly use this abstract state transition graph to define a skill hierarchy.
@@
Lastly, we note that the various characterisations of useful skills proposed in the literature, including the one proposed here, are not necessarily competitors. To solve complex tasks, it is likely that an agent will need to use many different types of skills. An important avenue for future work is exploring how different ideas on skills discovery can be used together to enable agents that can autonomously develop complex and varied skill hierarchies.
@@
We would like to thank the members of the Bath Reinforcement Learning Laboratory for their constructive feedback. This research was supported by the Engineering and Physical Sciences Research Council [EP/R513155/1] and the University of Bath. This research made use of Hex, the GPU Cloud in the Department of Computer Science at the University of Bath.